12 men can complete a work in 18 days. After 5 days 6 more men joined the work. In how many days will the

remaining work be completed?
Answer #1

Work done by 12 men in 5 days => (12 x 18)/1 

= (12 x 5)/x => x = 5/18

Remaining work = 13/18

Time taken by 12 + 6 men to do the remaining 

work => (12 x 18)/1 = (18 x X)/(13/18) => X = 26/3

Therefore time taken to complete the work is 8 2/3days.


answered by: Arjun Ananth